OpenStudy (anonymous):

I need help with 3 Algebra 1 problems . Medal & Fan !

OpenStudy (anonymous):

What problems do you need help with?

OpenStudy (anonymous):

The perimeter of a rectangle is 64 units. Can the length x of the rectangle can be 20 units when its width y is 11 units? No, the rectangle cannot have x = 20 and y = 11 because x + y ≠ 64. No, the rectangle cannot have x = 20 and y = 11 because x + y ≠ 32. Yes, the rectangle can have x = 20 and y = 11 because x + y is less than 64. Yes, the rectangle can have x = 20 and y = 11 because x + y is less than 32.

OpenStudy (anonymous):

The table shows the solution to the equation |2x – 3| - 1 = 2: Step 1 |2x – 3| = 2 + 1 Step 2 |2x – 3| = 3 Step 3 2x – 3 = 3 or 2x + 3 = 3 Step 4 2x = 6 or 2x = 0 Step 5 x = 3 or x = 0 Which is the first incorrect step? Step 1 Step 3 Step 5 Solution is correct

OpenStudy (anonymous):

Choose the missing step in the given solution to the inequality -x - 10 > 14 + 2x. -x - 10 > 14 + 2x -3x -10 > 14 x < -8 -3x > 24 -3x < 4 -3x > -4 8 < x

OpenStudy (anonymous):

ok so on the -x-10 > 14+2x whats missing is the 3rd step before you divide by the -3 you need to add youy 10 to the other side so that your variable is isolated. The missing step is -3x > 24 Does that make sense?

OpenStudy (anonymous):

Yea ! omg thank you . Can I ask you one more ?

OpenStudy (anonymous):

lol yea and ill answer the other 2 also

OpenStudy (anonymous):

On the \[\left| 2x-3 \right|=2+1\] Both answers are correct plug in both equations plugging in 3 gives l2*3-3l=3 l6-3l=3 3=3 on the x=0 l2*0-3l=3 l-3l=3 (because of the absolute values our - becomes + so 3=3 In conclusion none of the steps are wrong.

OpenStudy (anonymous):

If anything does not make sense just speak up

OpenStudy (anonymous):

On your first one about the perimeter of a rectangle. P=2l+2w basically adding up all the sides. If the perimeter did = 64 then 20+20+11+11would equal 64 and does it?

OpenStudy (anonymous):

looking at the answers they are asking just about x+y so only half the perimeter which would be half of 64. So does 20+11=32?

OpenStudy (anonymous):

Answer: No, the rectangle cannot have x=20 and y=11 because x+y does not equal 32

OpenStudy (anonymous):

Sorry i didnt see you replied ! other question was ; The sum of two consecutive odd integers is 244. What is the smaller integer? 119 125 123 121

OpenStudy (anonymous):

x+(x+2)=244 subtract the 2 from both sides 2x=242 divide by 2 to both sides x=121

OpenStudy (anonymous):

Thank you SO much !

OpenStudy (anonymous):

it was no problem :) is there any more I can help you with?

OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x - 1) -12? x -5 x -5 -5 x A number line with a closed circle on negative 5 and shading to the right

OpenStudy (anonymous):

ok im not sure what you are looking at so let me get this right. the equation is \[2(x-1)-12\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well i get x=7 is that what the equation is asking for?

OpenStudy (anonymous):

I think it's like the line with the closed or open circle . idk what it's called we can skip that question .

OpenStudy (anonymous):

Jasmine wants to use her savings of $1128 to buy video games and music CDs. The total price of the music CDs she bought was $72. The video games cost $43 each. What is the maximum number of video games that Jasmine can buy with her savings? 24 25 26 27

OpenStudy (anonymous):

oh maybe the equation is 2(x-1)>12??

OpenStudy (anonymous):

2(x-1) > <- ( Line under that ) -12

OpenStudy (anonymous):

\[2(x-1) \ge -12\]

OpenStudy (anonymous):

yes (:

OpenStudy (anonymous):

ok! now i got it lol

OpenStudy (anonymous):

Haha sorry i didnt realize that it didnt show those

OpenStudy (anonymous):

oh no its cool haha i see what they want now

OpenStudy (anonymous):

so the answer would be \[x \ge -5 \] do you see how i get that?

OpenStudy (anonymous):

yeah I got the same but i was unsure of it

OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 8x - (5x + 4) -31? x -9 -9 x A number line with a closed circle on -9 and shading to the left A number line with a closed circle on -9 and shading to the right

OpenStudy (anonymous):

 ok bad drawing but thats what your graph will look like. make sense? your x needs to be greater than negative 5 so shade the bigger numbers. and its a closed circle because its greater then and equal to.